# Can LLMs Simulate Human Behavioral Variability? A Case Study in the Phonemic Fluency Task

### v1

@@ Abstract
Large language models (LLMs) are increasingly explored as substitutes for human participants in cognitive tasks, but their ability to simulate human behavioral variability remains unclear. This study examines whether LLMs can approximate individual differences in the phonemic fluency task, where participants generate words beginning with a target letter. We evaluated 34 model configurations, varying prompt specificity, sampling temperature, and model type, and compared outputs to responses from 106 human participants. While some configurations, especially Claude 3.7 Sonnet, matched human averages and lexical preferences, none reproduced the scope of human variability. LLM outputs were consistently less diverse and structurally rigid, and LLM ensembles failed to increase diversity. Network analyses further revealed fundamental differences in retrieval structure between humans and models. These results highlight key limitations in using LLMs to simulate human cognition and behavior.
@@
Mengyang Qiu1††thanks: Corresponding author. Zoe Brisebois1 Siena Sun2 1Department of Psychology, Trent University, Canada 2Speech-Language Pathology Program, Saint Elizabeth University, United States {mengyangqiu, zoebrisebois}@trentu.ca ssun@steu.edu
1 Introduction
@@
This lack of variability also becomes evident when examining how LLMs perform in tasks designed to probe semantic or associative structures. In a semantic fluency task where participants named as many animals as possible, Wang et al. (2025) found that LLM-generated semantic networks were structurally different from those of humans. Compared to human networks, those of LLMs exhibited weaker local associations, poorer global integration, and greater rigidity in semantic organization. Similarly, Haim et al. (2025) examined word association networks in a STEM-related mindset task and found that GPT-3.5 produced networks that were notably sparser and less interconnected than those generated by human participants.
It is worth noting that Wang et al. (2025) recognized this limitation and attempted to address it by prompting LLMs to role-play 30 different occupations, effectively simulating multiple distinct agents. However, the resulting semantic networks still failed to match the flexibility and associative structure of human data. Crucially, real human participants, even those drawn from relatively homogeneous groups such as college students, consistently exhibit meaningful individual differences. These differences may reflect not only variation in the amount, type, and organization of semantic knowledge, but also differences in executive function, and strategy use. Of course, because LLMs are optimized to encode and reproduce stable semantic relationships, consistency across prompts may reflect a design feature. From this perspective, the internal coherence and semantic regularity that make LLMs so effective at language modeling may also constrain their ability to simulate the variability that characterizes human cognition.
@@
The goal of the present study is to explore whether LLMs are more capable of simulating inter-participant variability in a less semantically driven challenge–namely, the phonemic fluency task–or whether their outputs remain rigid, perhaps driven by frequency-based or alphabetic heuristics. Alongside this primary question, we examine several additional factors that may influence the diversity and human-likeness of LLM-generated responses:
1.
Participant context prompts: We systematically vary the type of information provided to the model, including demographic details (age, education) and performance-level information (number of correct responses), to assess whether real participant metadata supports more individualized outputs.
2.
Sampling temperature: Building on prior work that links temperature to creative or divergent generation (Peeperkorn et al., 2024), we test whether adjusting this parameter affects response diversity and whether higher temperatures introduce human-like variation or merely add noise.
3.
Model architecture and reasoning capacity: We compare general-purpose LLMs (e.g., GPT-4.1) to reasoning-oriented models (e.g., O3-mini) to examine whether models designed to engage in multi-step thinking produce more flexible or human-aligned behavior.
4.
Ensemble production: Finally, we investigate whether sampling outputs from multiple models, rather than relying on single model outputs, can better approximate the range of responses observed across human participants. While recent ensemble methods have primarily focused on improving task accuracy by using techniques like sampling and majority voting to converge on a correct answer (Li et al., 2024), our goal is the opposite: to explore whether ensembling can simulate the distributional variability seen in human behavior.
@@
In preview, our findings suggest a clear answer: LLMs are highly capable of producing fluent and correct answers in the phonemic fluency task, but none of them match the full scope of human behavioral variability. The remainder of the paper is organized as follows. Section 2 describes the human F fluency dataset, the full range of LLM configurations tested, and our approach to simulating participant-level outputs under different prompt and temperature conditions. Section 3 evaluates how well LLMs replicate human response counts and lexical diversity across different prompting setups and models. Section 4 presents item-level analyses, examining how linguistic features (e.g., word frequency, age of acquisition, phonological neighbors) predict production frequency in humans versus models. Section 5 turns to structural comparisons of word co-occurrence networks, analyzing clustering, path length, and representational similarity. Section 6 concludes with a discussion of implications for using LLMs as behavioral proxies and offers recommendations for improving variability in future simulations.
2 Method
2.1 Human data source
@@
We used verbal fluency data collected by Qiu and Johns (2021)111The verbal fluency data are publicly available on the Open Science Framework under the Creative Commons Attribution 4.0 International License (CC BY 4.0). We adhered to all applicable terms for reuse and attribution., who investigated noun- and verb-based semantic fluency across two experiments. In both experiments, participants first completed a phonemic fluency task using the letter F as a familiarization trial before proceeding to semantic tasks. A total of 106 native English speakers (age: ; education: years, ) were recruited via Amazon Mechanical Turk and completed the tasks online. Audio recordings were obtained for all participants’ responses to the F fluency task.
@@
Following the original study’s procedure, we used the Google Speech-to-Text API222https://cloud.google.com/speech-to-text to transcribe the recordings. Transcripts were then manually reviewed to identify and exclude errors, including repetitions, out-of-category responses, same-lemma variants (e.g., fish, fishing, fished), and counting behavior, where multiple number words (e.g., four, five, fourteen) were produced.
Participants produced an average of 17.27 words (), which slightly decreased to 16.89 words () after manual corrections. This small reduction indicates overall good adherence to task instructions. However, this average was lower than participants’ performance on the semantic fluency tasks in the same study, especially the noun-based conditions (animals, and living or nonliving things), where the mean number of responses exceeded 22. This pattern highlights the greater difficulty of phonemic fluency relative to semantic fluency.
2.2 LLM simulation procedure
@@
We evaluated a diverse set of LLMs from six major providers. For closed-source models, we used the official APIs from OpenAI333https://openai.com/api/, Anthropic444https://www.anthropic.com/api, Google555https://ai.google.dev/gemini-api/docs, and xAI666https://x.ai/api. For open-source models, we accessed models from Meta and Alibaba via the Together.ai API777https://www.together.ai. Each configuration was prompted to simulate all 106 human participants.888All LLM-generated outputs used in this study will be made publicly available under the Creative Commons Attribution 4.0 International License (CC BY 4.0).
To assess behavioral realism, we systematically varied prompt specificity and sampling temperature. A representative full prompt is shown in Figure 1. This version included the participant age, education, and number of correct responses. We also tested three reduced-information variants: a demographic-only prompt (no number of correct responses), a performance-only prompt (no demographics), and a no-information prompt. All prompts followed the same task format and instruction wording as the human study, and models were instructed to output one word per line.
In total, 34 distinct model configurations were evaluated. Appendix A provides detailed specifications for each model, including provider, temperature settings, and prompt type.
3 Participant-Level Analysis
3.1 Number of responses
Models from leading LLM providers were evaluated on their ability to simulate human performance in the letter F fluency task, with additional testing of the latest models from OpenAI (GPT-4.1) and Anthropic (Claude 3.7 Sonnet) using limited participant information. Human participants produced an average of 16.89 correct responses within the one-minute time constraint (). Most model configurations, when provided with full information or at least performance information (i.e., number of correct responses), generated similar response counts to humans (after removing repetitions and other errors), suggesting generally successful adherence to the instructed constraints.
When provided with only demographic information (age and education), GPT-4.1 generated significantly more words () and Claude 3.7 Sonnet showed similar overproduction ()–both more than doubling human output. This pattern was even more pronounced in the no-information condition, with GPT-4.1 generating higher counts () and Claude 3.7 Sonnet producing comparable results ().
Surprisingly, O4-mini, one of OpenAI’s latest reasoning models, demonstrated extreme overestimation with demographic information ()–nearly 60 more words than human participants. Even when performance information was explicitly provided to this model, it still overproduced with considerable variance (), indicating inconsistent adherence to the performance constraints specified in the prompt. Interestingly, with the full prompt that included both demographic and performance information, O4-mini accurately reproduced human-like performance in terms of response count.
For several other models, even with full participant information, we observed unstable production patterns characterized by extreme bursts of responses, sometimes generating around 100 words for some participants while appropriately following instructions for others. In particular, OpenAI’s GPT-4-turbo () and O3 () models generated excessive output with high variance, suggesting these models inconsistently applied the performance constraints specified in the prompt.
Table 1 summarizes the 11 model configurations that significantly overestimated human performance at the group level. Additionally, four more configurations, including Google’s Gemini 2.0 and Gemini 2.5 (with and without thinking enabled), and Alibaba’s Qwen-2.5-72B, showed similar bursts of responses for individual participants, though these did not reach statistical significance at the group level. These 15 model configurations were excluded from subsequent analyses to focus on models with more realistic human simulation.
Provider | Model | Prompt Type | Mean SD | Diff. | t |
Human | – | – | – | – | |
Models with Demographic-Info or No-Info Prompts | |||||
OpenAI | GPT-4.1 | Demographic | |||
No | |||||
O4-mini (reasoning) | Demographic | ||||
[…]
No | |||||
Models with Performance-Info or Full Prompts | |||||
OpenAI | GPT-4-turbo | Full | |||
O3 (reasoning) | Full | ||||
O4-mini (reasoning) | Performance | ||||
[…]
Llama 3.3 70B | Full | ||||
Note: Diff. = mean difference from human performance; ; |
3.2 Variability of responses
@@
To evaluate the variability of responses across human participants and LLM-simulated participants, we computed the Type-to-Token Ratio (TTR) and the Idiosyncratic Type-to-Total Type Ratio (ITTR). TTR captures lexical diversity by dividing the number of unique words (types) by the total number of responses (tokens). A higher TTR indicates a wider range of vocabulary produced. ITTR measures how many of those types were idiosyncratic (i.e., produced by only one participant) relative to the total number of types. This provides insight into how consistent or individualized the generated words are across participants within the group (Castro et al., 2021).
@@
As shown in Table 2, human participants produced the highest number of unique types (476) and idiosyncratic types (201), resulting in a TTR of 0.27 and an ITTR of 0.42. No LLMs approached this level of variability. The most diverse LLM output came from Claude 3.7 Sonnet at temperature 1.0 (226 types), followed closely by O3-mini, the best-performing reasoning model (184 types). However, their ITTRs (0.32 and 0.28, respectively) remained well below the human benchmark.
As expected, increasing temperature led to more diverse responses for both GPT-4.1 and Claude 3.7 Sonnet, with higher TTR and ITTR values at higher temperatures, though they still failed to match human-like variability. We also tested how prompt specificity affected variability: for GPT-4.1, removing demographic details while retaining performance information slightly increased TTR and ITTR, but for Claude 3.7, the pattern reversed. A trend across model generations also emerged: holding temperature constant at 0.7, lexical diversity decreased from GPT-3.5 to GPT-4o to GPT-4.1. This suggests that newer models may prioritize consistency and safety over variability.
Finally, we examined whether ensembles of LLMs could recover some of the lost variability by combining outputs from diverse models. Three ensemble strategies were tested: sampling from the top five models in terms of type count, from all models with at least 100 types (9 in total), and from all models that received the full prompt (17 in total). As shown in the bottom portion of Table 2, none of the ensemble approaches exceeded the lexical diversity of the best individual model. The top-five ensemble achieved 225 types, nearly identical to Claude 3.7 Sonnet at temperature 1.0. This suggests that between-model differences in lexical choices are limited, and ensemble generation alone is insufficient to reach human-level diversity.
Provider | Model | Temperature | Types | Idio. Types | Tokens | TTR | ITTR |
Human | – | 476 | 201 | 1790 | 0.27 | 0.42 | |
OpenAI | |||||||
GPT-4.1 | 0.3 | 57 | 10 | 1766 | 0.03 | 0.18 | |
0.7 (default) | 69 | 18 | 1778 | 0.04 | 0.26 | ||
1.1 | 108 | 37 | 1783 | 0.06 | 0.34 | ||
1.5 | 138 | 38 | 1777 | 0.08 | 0.28 | ||
[…]
GPT-4o | 0.7 (default) | 128 | 38 | 1773 | 0.07 | 0.30 | |
GPT-3.5 | 0.7 (default) | 148 | 48 | 1783 | 0.08 | 0.32 | |
[…]
O4-mini (reasoning) | – | 88 | 24 | 1790 | 0.05 | 0.27 | |
Anthropic | |||||||
Claude 3.7 Sonnet | 0.3 | 97 | 28 | 1787 | 0.05 | 0.29 | |
0.7 | 150 | 39 | 1788 | 0.08 | 0.26 | ||
1.0 | 226 | 73 | 1781 | 0.13 | 0.32 | ||
[…]
Claude 3.7 Sonnet
0.7 | 137 | 35 | 1787 | 0.08 | 0.26 |
Claude 3.7 Sonnet
(thinking enabled) |
– | 137 | 43 | 1788 | 0.08 | 0.31 | |
xAI | |||||||
Grok-3 | 0.7 | 100 | 22 | 1790 | 0.06 | 0.22 | |
1.0 | 94 | 18 | 1790 | 0.05 | 0.19 | ||
[…]
1.0 | 71 | 19 | 1793 | 0.04 | 0.27 | ||
Open-Source (Meta) | |||||||
Llama 4 Scout | default | 83 | 25 | 1783 | 0.05 | 0.30 | |
LLM Ensembles | |||||||
Top 5 Models
(Word Types) |
– | 225 | 75 | 1778 | 0.13 | 0.33 | |
100+ Word Types
(9 models) |
– | 180 | 46 | 1785 | 0.10 | 0.26 | |
All-Model Mix
(17 models) |
– | 169 | 45 | 1780 | 0.09 | 0.27 |
4 Item-Level Analysis
4.1 Distribution of production frequency

### v2

@@ Abstract
Large language models (LLMs) are increasingly explored as substitutes for human participants in cognitive tasks, but their ability to simulate human behavioral variability remains unclear. This study examines whether LLMs can approximate individual differences in the phonemic fluency task, where participants generate words beginning with a target letter. We evaluated 34 distinct models across 45 configurations from major closed-source and open-source providers, and compared outputs to responses from 106 human participants. While some models, especially Claude 3.7 Sonnet, approximated human averages and lexical preferences, none reproduced the scope of human variability. LLM outputs were consistently less diverse, with newer models and thinking-enabled modes often reducing rather than increasing variability. Network analysis further revealed fundamental differences in retrieval structure between humans and the most human-like model. Ensemble simulations combining outputs from diverse models also failed to recover human-level diversity, likely due to high vocabulary overlap across models. These results highlight key limitations in using LLMs to simulate human cognition and behavior.
Keywords: large language models, phonemic fluency, behavioral variability, cognitive simulation
@@
Abstract content
1. Introduction
@@
This lack of variability is also evident when examining how LLMs perform in tasks designed to probe semantic or associative structures. In a semantic fluency task where participants named as many animals as possible, Wang et al. (2025b) found that LLM-generated semantic networks were structurally different from those of humans. Compared to human networks, those of LLMs exhibited weaker local associations, poorer global integration, and greater rigidity in semantic organization. Similarly, Haim et al. (2025) examined word association networks in a STEM-related mindset task and found that GPT-3.5 produced networks that were notably sparser and less interconnected than those generated by human participants.
It is worth noting that Wang et al. (2025b) recognized this limitation and attempted to address it by prompting LLMs to role-play 30 different occupations, effectively simulating multiple distinct participants. However, the resulting semantic networks still failed to match the flexibility and associative structure of human data. Crucially, real human participants, even those drawn from relatively homogeneous groups such as college students, consistently exhibit meaningful individual differences. These differences may reflect not only variation in the amount, type, and organization of semantic knowledge, but also differences in executive function, and strategy use. Of course, because LLMs are optimized to encode and reproduce stable semantic relationships, consistency across prompts may reflect a design feature. From this perspective, the internal coherence and semantic regularity that make LLMs so effective at language modeling may also constrain their ability to simulate the variability that characterizes human cognition.
@@
The goal of the present study is to explore whether LLMs can simulate inter-participant variability in phonemic fluency, or whether their outputs remain rigid, perhaps driven by frequency-based or alphabetic heuristics. Unlike previous work that introduced arbitrary personas such as fictional occupations (Wang et al., 2025b), we ground our simulations in real participant metadata. Extensive research in psychology has consistently shown that age and education are among the most reliable predictors of verbal fluency performance (Tombaugh et al., 1999; Gladsjo et al., 1999; Villalobos et al., 2023). By providing LLMs with demographic and performance information drawn directly from real human participants, the simulation can be more meaningfully compared against actual human data.
Furthermore, previous studies comparing LLM and human cognitive and language performance have typically relied on one or a small number of models. The present study offers a substantially more comprehensive evaluation, testing distinct models across configurations from major closed-source and open-source providers. This broad coverage allows us to identify patterns that hold across model families and architectures, rather than drawing conclusions from a single system.
It also enables us to address a fundamental concern about LLM-based simulation: using the same model to simulate multiple participants means those simulated individuals inherently share the same foundational knowledge, making it difficult to generate authentically varied outputs regardless of prompt manipulation (Wang et al., 2025a). To overcome this limitation, we investigate whether combining outputs from diverse models through ensemble sampling—where each simulated participant’s response is drawn from a randomly selected model—can better approximate human-level diversity, leveraging architectural and training differences across models to collectively recover variability that no single model captures alone.
@@
In preview, our findings suggest a clear answer: LLMs are highly capable of producing fluent and correct responses in the phonemic fluency task, but none of them—individually or in ensemble—match the full scope of human behavioral variability. The remainder of the paper is organized as follows. Section 2 describes the human F fluency dataset and the LLM simulation procedure. Section 3 evaluates participant-level response counts and lexical diversity. Section 4 presents item-level analyses of production frequency distributions and linguistic predictors. Section 5 compares word co-occurrence network structures. Section 6 examines ensemble-based diversity. Section 7 concludes with a discussion of implications and future directions.
2. Method
2.1. Human data source
@@
We used phonemic fluency data collected by Qiu and Johns (2021)111The verbal fluency data are publicly available on the Open Science Framework under the Creative Commons Attribution 4.0 International License (CC BY 4.0)., who investigated noun- and verb-based semantic fluency across two experiments (see also Qiu et al., 2021). In both experiments, participants first completed a phonemic fluency task using the letter F as a familiarization trial before proceeding to semantic tasks. A total of 106 native English speakers (age: ; education: years, ) were recruited via Amazon Mechanical Turk and completed the tasks online. Participants were instructed to produce as many words as possible beginning with the letter F within one minute, excluding proper nouns, numbers, and same-lemma variants. Audio recordings were collected for all participants.
@@
Following the original study’s procedure, we used the Google Speech-to-Text API222https://cloud.google.com/speech-to-text to transcribe the recordings, followed by manual review to exclude errors such as repetitions, out-of-category responses, same-lemma variants (e.g., fish, fishing, fished), and counting behavior. After corrections, participants produced an average of 16.89 words ().
2.2. LLM simulation procedure
@@
We evaluated a diverse set of closed-source and open-source LLMs. For closed-source models, we used the official APIs from OpenAI333https://openai.com/api/, Anthropic444https://www.anthropic.com/api, Google555https://ai.google.dev/gemini-api/docs, and xAI666https://x.ai/api. For open-source models, we accessed models from Meta, Moonshot, Alibaba, Zhipu AI, DeepSeek, and Mistral via the Together.ai API777https://www.together.ai.
All models received the same full prompt, which included the participant’s age, education level, and number of correct responses, along with task instructions identical to those given to human participants (see Appendix A for the complete prompt). The prompt instructed models to role-play as the corresponding human participant and generate words beginning with F under a simulated one-minute constraint, outputting one word per line with no additional text. Each model was prompted independently 106 times, once per human participant, with each prompt constituting a separate API call with no memory of previous calls.
In total, we evaluated 34 distinct models. Given that many recent models support hybrid thinking or reasoning modes, we additionally tested these models with and without extended thinking enabled, and at different reasoning effort levels where applicable, resulting in 45 configurations.
3. Participant-Level Analysis
3.1. Number of responses
Human participants produced an average of 16.89 correct responses within the one-minute time constraint (). To evaluate whether LLMs adhered to the performance constraints specified in the prompt, we computed the Mean Absolute Error (MAE) between LLM-generated and human response counts for each configuration. Configurations with (i.e., within 10% of the human mean) were considered to have successfully simulated human-like production rates.
Of the 45 configurations tested, 33 met this criterion, indicating that the majority of models generally adhered to the instructed constraints. The remaining 12 configurations significantly overproduced relative to human participants. Among these, OpenAI’s GPT-4 Turbo () and o3 () showed the most extreme overproduction, generating outputs several times longer than human responses. Surprisingly, two of Anthropic’s latest models, Claude 4.6 Sonnet (Thinking) () and Claude 4.6 Opus (), also failed to adhere to the performance constraints despite the explicit inclusion of the number of correct responses in the prompt.
Among the 33 configurations with , we further examined whether any individual simulations deviated substantially from the corresponding human participant’s response count. Configurations containing any simulation where were flagged as having outliers. Of the 33 configurations, 21 produced no outliers across all 106 simulations and were retained for subsequent analyses. Details of included and excluded configurations are provided in Appendix B.
3.2. Variability of responses
@@
To evaluate the variability of responses across human participants and LLM-simulated participants, we computed the Type-to-Token Ratio (TTR) and the Idiosyncratic Type-to-Total Type Ratio (ITTTR). TTR captures lexical diversity by dividing the number of unique words (types) by the total number of responses (tokens). A higher TTR indicates a wider range of vocabulary produced. ITTTR measures how many of those types were idiosyncratic (i.e., produced by only one participant) relative to the total number of types. This provides insight into how consistent or individualized the generated words are across participants within the group (Castro et al., 2021).
@@
As shown in Table 1, human participants produced the highest number of unique types () and idiosyncratic types (), resulting in a TTR of and an ITTTR of . No LLM approached this level of variability. The most diverse LLM output came from Claude 3.7 Sonnet ( types, TTR , ITTTR ), followed by o3-mini ( types, TTR , ITTTR ). Even these best-performing models produced fewer than half the unique types observed in human data.
Several patterns emerged across model families. For Anthropic models, non-thinking configurations consistently produced more diverse outputs than their thinking-enabled counterparts (e.g., Claude 3.7 Sonnet: vs. types; Claude 4.5 Sonnet: vs. types; Claude 4 Sonnet: vs. types). Additionally, the earlier Claude 3.7 Sonnet substantially outperformed more recent Anthropic models in lexical diversity, suggesting that newer models may prioritize consistency over variability.
For OpenAI models, reasoning capacity appeared to benefit diversity. GPT-5.2 showed a clear gradient, with higher reasoning effort producing more unique types (High: , Medium: , Low: ). The most diverse OpenAI model, o3-mini ( types), is itself a reasoning-focused model. Despite these differences across models, all LLM configurations remained well below human-level variability in both TTR and ITTTR, highlighting a persistent tendency toward output rigidity across architectures and providers.
| Model | Types | Idio. Types | Tokens | TTR | ITTTR |
| Human | 476 | 201 | 1790 | 0.27 | 0.42 |
| OpenAI | |||||
| o3-mini | 193 | 58 | 1790 | 0.11 | 0.30 |
| GPT-5.2 (High Reasoning) | 158 | 40 | 1790 | 0.09 | 0.25 |
| GPT-5.2 (Medium Reasoning) | 149 | 38 | 1790 | 0.08 | 0.26 |
| GPT-3.5 Turbo | 148 | 48 | 1783 | 0.08 | 0.32 |
| GPT-5.2 (Low Reasoning) | 139 | 35 | 1790 | 0.08 | 0.25 |
| GPT-5 | 130 | 33 | 1790 | 0.07 | 0.25 |
| GPT-4o | 128 | 38 | 1773 | 0.07 | 0.30 |
| GPT-5 mini | 95 | 18 | 1790 | 0.05 | 0.19 |
| GPT-4.1 | 73 | 20 | 1790 | 0.04 | 0.27 |
| Anthropic | |||||
| Claude 3.7 Sonnet | 226 | 73 | 1781 | 0.13 | 0.32 |
| Claude 3.7 Sonnet (Thinking) | 137 | 43 | 1788 | 0.08 | 0.31 |
| Claude 4.5 Sonnet | 110 | 32 | 1784 | 0.06 | 0.29 |
| Claude 4.5 Sonnet (Thinking) | 77 | 18 | 1790 | 0.04 | 0.23 |
| Claude 4 Sonnet | 74 | 12 | 1795 | 0.04 | 0.16 |
| Claude 4 Sonnet (Thinking) | 73 | 18 | 1789 | 0.04 | 0.25 |
| Gemini 2.5 Flash | 149 | 51 | 1794 | 0.08 | 0.34 |
| Gemini 2.5 Pro | 83 | 20 | 1790 | 0.05 | 0.24 |
| Gemini 3 Pro | 73 | 20 | 1790 | 0.04 | 0.27 |
| xAI | |||||
| Grok-4.1-fast (Reasoning) | 62 | 8 | 1787 | 0.03 | 0.13 |
| Open-Source Models | |||||
| Kimi K2 | 96 | 25 | 1787 | 0.05 | 0.26 |
| Qwen3-2507 | 93 | 22 | 1781 | 0.05 | 0.24 |
4. Item-Level Analysis
4.1. Distribution of production frequency